if Cotangent theta is 1/2 and sin <0. what are the other trig functions?
Answer by stanbon(75887)

if Cot(t) = 1/2 and sin < 0. what are the other trig functions?
----
Assuming t is in QI,
Since cot = x/y, x = 1 and y = 2
Then r = sqrt[1^2+2^2] = sqrt(5)
-----
sin = y/r = 2/sqrt(5)
cos = x/r = 1/sqrt(5)
tan = y/x = 2
-----
csc = r/y = sqrt(5)/2
sec = r/x = sqrt(5)
cot = x/y = 1/2
------------------
Cheers,
Stan H.
------------------